# Lightweight Deep Learning-Based Channel Estimation for RIS-Aided Extremely Large-Scale MIMO Systems on Resource-Limited Edge Devices

### v1

@@ Abstract
Next-generation wireless technologies such as 6G aim to meet demanding requirements such as ultra-high data rates, low latency, and enhanced connectivity. Extremely Large-Scale MIMO (XL-MIMO) and Reconfigurable Intelligent Surface (RIS) are key enablers, with XL-MIMO boosting spectral and energy efficiency through numerous antennas, and RIS offering dynamic control over the wireless environment via passive reflective elements. However, realizing their full potential depends on accurate Channel State Information (CSI). Recent advances in deep learning have facilitated efficient cascaded channel estimation. However, the scalability and practical deployment of existing estimation models in XL-MIMO systems remain limited. The growing number of antennas and RIS elements introduces a significant barrier to real-time and efficient channel estimation, drastically increasing data volume, escalating computational complexity, requiring advanced hardware, and resulting in substantial energy consumption. To address these challenges, we propose a lightweight deep learning framework for efficient cascaded channel estimation in XL-MIMO systems, designed to minimize computational complexity and make it suitable for deployment on resource-constrained edge devices. Using spatial correlations in the channel, we introduce a patch-based training mechanism that reduces the dimensionality of input to patch-level representations while preserving essential information, allowing scalable training for large-scale systems. Simulation results under diverse conditions demonstrate that our framework significantly improves estimation accuracy and reduces computational complexity, regardless of the increasing number of antennas and RIS elements in XL-MIMO systems.
Index Terms:
Reconfigurable Intelligent Surface, B6G, Deep Learning, UNet 3+, Channel Estimation, Extremely Large-Scale MIMO, Spatial Correlation in Wireless Channels.
@@ I Introduction
Next-generation communication technologies, such as 6G and beyond, are expected to meet the rapidly growing demand for advanced service requirements [1]. These include ultra-high data rates, extremely low latency, minimal bit error rates, enhanced device connectivity, and much more. To address these challenges, Extremely Large-Scale MIMO (XL-MIMO) has emerged as a transformative paradigm, offering significantly higher spectral and energy efficiency while ensuring reliable communication [2]. Unlike conventional MIMO systems, XL-MIMO integrates thousands of antennas in a compact space, fundamentally altering electromagnetic characteristics [3]. This not only expands the number of antennas, but also enhances spectral efficiency and spatial degrees of freedom, enabling unprecedented improvements in wireless communication performance [4]. Building on existing advantages, the integration of Reconfigurable Intelligent Surfaces (RIS) with XL-MIMO presents a promising approach to address the growing demands of 6G and beyond networks [1].
RIS technology is emerging as a transformative enabler, offering the ability to dynamically shape wireless channels between users and the Base Station (BS), thereby enhancing the overall performance of the system [5]. RIS panels consist of tunable electromagnetic elements, each independently controlled to direct incident waves towards targeted directions [6, 7]. This adaptability optimizes the propagation environment, significantly boosting the spectral efficiency [8]. Furthermore, RIS elements are energy efficient and passive, seamlessly integrated into surfaces such as walls and ceilings, making them a low-energy, cost-effective solution for significantly improving signal quality and network coverage [9].
@@
Realizing the performance gains of RIS-enabled XL-MIMO systems in practice requires optimizing the design parameters, which are highly dependent on accurate Channel State Information (CSI) [10]. However, obtaining CSI in RIS is challenging because RIS elements are passive and do not receive or transmit signals directly [11]. Since individual channels between user–RIS and RIS–BS cannot be estimated separately, only the cascaded channel can be inferred, which encompasses the link from the user to the BS via the RIS. This limitation presents distinct challenges in accurately estimating the overall channel. Firstly, unlike traditional Rayleigh fading channels, RIS-based channels deviate from conventional models, resulting in a performance gap between Minimum Mean Square Error (MMSE) and Linear-MMSE (LMMSE) estimators [1]. Moreover, both RIS and antenna arrays consist of numerous elements, which complicates cascaded channel estimation and significantly expands the parameter space, making channel estimation resource-intensive for edge devices [12, 6]. Consequently, these challenges limit system performance and hinder the large-scale deployment of RIS-assisted XL-MIMO systems. This paper proposes a power-efficient data-driven methodology for cascaded channel estimation at the edge devices for RIS-enabled XL-MIMO systems.
@@
Recent work in cascaded channel estimation can be broadly categorized into two main sets of approaches: conventional model-driven methods [6, 9, 10] and data-driven techniques [11, 12, 13, 14, 15, 16]. Conventional model-driven methods frame cascaded channel estimation as an optimization problem, solved iteratively to improve accuracy. Although these methods have achieved notable advancements, the need for multiple iterative steps in real-time scenarios makes them time-consuming, which is impractical for future low-latency communications [11]. In contrast, data-driven approaches (such as machine and deep learning) are gaining popularity. These techniques offer minimal inference time, making them well suited for real-world, time-sensitive scenarios. Although data-driven methods require relevant training data, they can significantly enhance the efficiency and accuracy of cascaded channel estimation once trained [15].
@@
Recent advances in deep learning have substantially contributed to RIS-based communication systems, delivering robust performance across diverse environmental conditions and enabling low-complexity inference. Various deep learning approaches have been proposed to improve cascaded channel estimation for massive MIMO systems, significantly improving overall system performance [15, 11, 12, 13, 14, 16]. However, these techniques face scalability challenges and are typically limited to massive MIMO systems, where only tens of antennas and BS elements are considered. Extending the training setup to RIS-enabled XL-MIMO systems, involving hundreds of RIS elements and thousands of antennas as proposed in these models, would not be a viable solution due to various reasons. Firstly, these techniques use tens of thousands of samples to train their models. Processing and storing such large datasets in RIS-assisted XL-MIMO-based systems can lead to inefficiencies, introduce substantial computational and memory demands, and increase both training time and resource consumption [11]. Furthermore, training these models on large datasets requires substantial energy resources, making them not only energy inefficient but also raising environmental concerns due to increasing energy consumption. Additionally, the need for highly specialized high-performance hardware to handle the training of such expensive systems adds considerable costs, making it financially challenging to scale these solutions. Finally, with edge computing gaining attention in recent years as a key enabler for distributed low-latency applications, the deployment of these complex models in edge environments becomes a significant obstacle [17]. The limited computational resources and energy constraints at the edge create a major barrier, hindering the practical implementation of these techniques in future communication networks.
@@
Addressing the aforementioned challenges requires a compact and efficient representation of the cascaded channel. We therefore exploit the spatial correlation present in the cascaded channel, as discussed in [18, 19, 20], to enable computationally efficient model training suitable for edge devices in resource-limited settings. To take advantage of this property, we introduce a patch-based training framework that leverages the inherent spatial correlations in the cascaded channel. In this framework, small patches are extracted from each data point to encapsulate the essential features required for accurate channel estimation. This patch-wise processing significantly reduces input dimensionality and overall system complexity, thereby enabling scalable and computationally efficient training, particularly suitable for XL-MIMO systems deployed at the edge. Furthermore, employing efficient deep learning techniques, such as autoencoders, can transform these representative patches into compact latent representations, thus facilitating more efficient learning of spatial dependencies [21]. These latent features effectively preserve and highlight spatial dependencies between RIS elements and BS antennas, facilitating the learning of inherent correlations while substantially reducing overall training complexity. Furthermore, leveraging autoencoders also helps break the correlation of noise in the data, effectively isolating meaningful patterns from unwanted distortions [22].
@@
Emerging autoencoder models, such as UNet, have shown remarkable success in noisy data segmentation and reconstruction by capturing underlying data structures, even in data-limited settings [21]. Unlike conventional autoencoders, UNet enables more precise feature extraction and better performance in complex tasks such as channel denoising. For example, the UNet model has demonstrated strong potential for channel estimation in Massive MIMO systems [23]. Extending the capabilities of the UNet, the UNet3+ architecture leverages multi-scale feature aggregation to better extract relevant structures from noisy inputs, improving denoising performance [24]. Therefore, we adopt the UNet 3+ architecture in our proposed channel estimation framework, leveraging its ability to effectively capture spatial features and improve estimation accuracy, especially in resource-constrained environments such as edge devices. Overall, this is the first work to propose an efficient solution for channel estimation in RIS-assisted XL-MIMO systems. Moreover, we formulate the channel estimation task as a denoising problem, where the goal is to recover clean channel information from noisy observations. The proposed method significantly improves the accuracy of the estimation and reduces computational complexity, regardless of the increasing number of antennas and RIS elements in XL-MIMO systems. Extensive simulations under diverse environmental conditions validate the scalability and effectiveness of the proposed framework, demonstrating substantial gains in cascaded channel estimation performance and overall system efficiency for RIS enabled XL-MIMO deployments.
@@ I-A Literature Review
Efficient cascaded channel estimation has attracted significant attention in recent years due to the promising benefits of RIS. Existing research efforts can be broadly categorized into two main approaches: traditional model-driven methods and data-driven techniques. Traditional model-based approaches for cascaded channel estimation remain widely used due to their ability to operate without training data, allowing direct deployment without prior model training. In [6], the authors tackled the cascaded channel estimation problem by exploiting the sparsity and structural correlations in mmWave environments. Their primary optimizations focused on adjusting the phase shifts of RIS elements and minimizing the pilot overhead required for successive channel updates. The simulation results confirmed the effectiveness of their approach, demonstrating a substantial reduction in training overhead. Moreover, in [7], a binary reflection-based channel estimation approach is proposed, wherein a single element is activated at a time while others remain off, and channel estimates are obtained using the least squares method.
@@
Furthermore, in [9], the authors formulated the cascaded channel estimation task using an alternating optimization algorithm combined with a covariance method, which they solved through a successive convex approximation algorithm. Their findings demonstrated enhanced channel estimation accuracy while reducing pilot overhead. The authors in [10] proposed a channel estimation protocol that leveraged compressive sensing techniques to estimate the cascaded channel. Their approach reduced training overhead by formulating the estimation task as a sparse recovery problem. They introduced a two-step joint channel estimation framework to further improve the estimation performance that exploited the block-sparsity structure shared among multiple users. Despite the adequate performance of these approaches  [6, 9, 10], they rely on numerous iterative steps to achieve optimal results, which limits their efficiency in real-time, low-latency applications when applied to XL-MIMO systems involving hundreds of RIS elements and thousands of antennas. As a result, model-based approaches become impractical for such large-scale deployments due to their excessive complexity and resource demands.
@@
The pioneering work that introduced deep learning for cascaded channel estimation was presented in [15]. The proposed framework consisted of two convolutional neural networks (CNNs), each comprising nine layers, designed to estimate both the direct channel between the BS and the user and the cascaded channel between the BS and the user through an RIS. The simulation results validated the effectiveness of the proposed model, demonstrating improvements over traditional methods such as Least Squares (LS) and multilayer perceptron. Moreover, in [11], the authors proposed a deep residual network (CDRN) method using a CNN-based architecture to effectively remove residual noise from noisy pilot signals, ultimately enhancing estimation accuracy. Additionally, [12] introduced a bidirectional long short-term memory (Bi-LSTM) approach for channel estimation. Bi-LSTM processed past and future information simultaneously through two LSTM layers, one forward and one backward, thereby enhancing feature extraction and achieving superior accuracy in channel estimation compared to other recurrent neural networks.
@@
Similarly, the work in [13] interprets the received pilot signal as a low-resolution image and formulates channel estimation as a super-resolution reconstruction problem. This approach employed a CNN to extract detailed features, followed by a denoising CNN to eliminate additive noise. The simulation results validated the effectiveness of this scheme. Another study in [16] proposed a model-driven deep unfolding approach for cascaded channel estimation. The proposed scheme leverages the rank deficiency of the cascaded channel to achieve accurate estimation while significantly reducing communication overhead. In [14], a self-supervised learning approach was proposed for channel estimation without the need for a ground truth table. This method generated inputs by adding noise to the received signal, while the noisy signal served as the data label. Although this approach yielded notable results, it faced limitations in real-world scenarios where accurate noise variance estimates were often unavailable. Although the techniques mentioned above [15, 11, 12, 13, 14, 16] demonstrate satisfactory performance, they still encounter challenges related to scalability and estimation accuracy, and rely on unrealistic dataset assumptions that may not adequately reflect real-world scenarios.
@@
A significant limitation of the data-driven techniques, as mentioned earlier, was that most proposed schemes were trained and evaluated on small-scale systems, typically involving only tens of RIS and antenna elements. Although these schemes perform effectively in moderate-scale scenarios, their scalability becomes a critical limitation when applied to large-scale systems involving hundreds of RIS elements and thousands of antennas, rendering them impractical for RIS-enabled XL-MIMO deployments [5]. Training such large systems requires complex and high-end infrastructure to handle the computational demands, posing a significant challenge to the feasibility of these approaches. However, several studies [5, 6, 10] have highlighted the presence of spatial correlation in the cascaded channel. In particular, some works have conducted a detailed analysis of spatially correlated RIS with antenna elements, providing valuable insights into the impact of correlation on system performance [18, 19, 20]. Therefore, this paper leverages the spatial correlation inherent in the cascaded channel to reduce model complexity, enabling efficient operation in XL-MIMO systems on resource-constrained devices. To effectively train our model, we employ an autoencoder-based architecture. The work presented in [23] has used an autoencoder-based architecture (UNet) combined with ResNet to capture spatial correlations between inputs and outputs, achieving strong performance in mmWave massive MIMO systems.

### v2

@@ Abstract
Next-generation wireless technologies such as Sixth Generation (6G) aim to meet demanding requirements such as ultra-high data rates, low latency, and enhanced connectivity. Extremely Large-Scale MIMO (XL-MIMO) and Reconfigurable Intelligent Surface (RIS) are key enablers, with XL-MIMO boosting spectral and energy efficiency through numerous antennas, and RIS offering dynamic control over the wireless environment via passive reflective elements. However, realizing their full potential depends on accurate Channel State Information (CSI). Recent advances in deep learning have facilitated efficient cascaded channel estimation. However, the scalability and practical deployment of existing estimation models in XL-MIMO systems remain limited. The growing number of antennas and RIS elements introduces a significant barrier to real-time and efficient channel estimation, drastically increasing data volume, escalating computational complexity, requiring advanced hardware, and resulting in substantial energy consumption. To address these challenges, we propose a lightweight deep learning framework for efficient cascaded channel estimation in XL-MIMO systems, designed to minimize computational complexity and make it suitable for deployment on resource-constrained edge devices. Using spatial correlations in the channel, we introduce a patch-based training mechanism that reduces the dimensionality of input to patch-level representations while preserving essential information, allowing scalable training for large-scale systems. Simulation results under diverse conditions demonstrate that our framework significantly improves estimation accuracy and reduces computational complexity, thereby maintaining robust performance even with increasing system dimensions in XL-MIMO networks.
@@ I Introduction
Next-generation communication technologies, such as Sixth Generation (6G) and beyond, are expected to meet the rapidly growing demand for advanced service requirements [RIS6G]. These include ultra-high data rates, extremely low latency, minimal bit error rates, enhanced device connectivity, and much more. To address these challenges, Extremely Large-Scale MIMO (XL-MIMO) has emerged as a transformative paradigm, offering significantly higher spectral and energy efficiency while ensuring reliable communication [XL1]. Unlike conventional MIMO systems, XL-MIMO integrates thousands of antennas in a compact space, fundamentally altering electromagnetic characteristics [XL2]. This not only expands the number of antennas, but also enhances spectral efficiency and spatial degrees of freedom, enabling unprecedented improvements in wireless communication performance [XL3]. Building on existing advantages, the integration of Reconfigurable Intelligent Surfaces (RIS) with XL-MIMO presents a promising approach to address the growing demands of 6G and beyond networks [RIS6G]. RIS technology is emerging as a transformative enabler, offering the ability to dynamically shape wireless channels between users and the Base Station (BS), thereby enhancing the overall performance of the system [RIS]. RIS panels consist of tunable electromagnetic elements, each independently controlled to direct incident waves towards targeted directions [OP1, BLMMSE]. This adaptability optimizes the propagation environment, significantly boosting the spectral efficiency [RIS1]. Furthermore, RIS elements are passive, energy-efficient, and easily integrated into surfaces, offering a low-cost solution to enhance signal quality and network coverage [OP2].
@@
Realizing the performance gains of RIS-enabled XL-MIMO systems in practice requires optimizing the design parameters, which are highly dependent on accurate Channel State Information (CSI) [OP3]. However, obtaining CSI in RIS is challenging because RIS elements are passive and do not receive or transmit signals directly [DL1]. Since individual channels between user–RIS and RIS–BS cannot be estimated separately, only the cascaded channel can be inferred, which encompasses the link from the user to the BS via the RIS. This limitation presents distinct challenges in accurately estimating the overall channel. Firstly, unlike traditional Rayleigh fading channels, RIS-based channels deviate from conventional models, resulting in a performance gap between Minimum Mean Square Error (MMSE) and Linear-MMSE (LMMSE) estimators [RIS6G]. Moreover, the large number of elements in RIS and antenna arrays complicates cascaded channel estimation, significantly expanding the parameter space and making it resource-intensive for edge devices [DL2, OP1]. Consequently, these challenges limit system performance and hinder the large-scale deployment of RIS-assisted XL-MIMO systems. This paper proposes a power-efficient data-driven approach for cascaded channel estimation at the edge devices for RIS-enabled XL-MIMO systems.
@@
Recent work in cascaded channel estimation can be broadly categorized into two main sets of approaches: conventional model-driven methods [OP1, OP2, OP3] and data-driven techniques [DL1, DL2, DL3, DL4, DL5, DL6]. Conventional model-driven methods frame cascaded channel estimation as an optimization problem, solved iteratively to improve accuracy. Although these methods have achieved notable advancements, the need for multiple iterative steps in real-time scenarios makes them time-consuming, which is impractical for future low-latency communications [DL1]. In contrast, data-driven approaches (such as machine and deep learning) are gaining popularity. These techniques offer minimal inference time, making them well suited for real-world, time-sensitive scenarios. Although data-driven methods require relevant training data, they can significantly enhance the efficiency and accuracy of channel estimation once trained [DL5].
@@
Recent advances in deep learning have substantially contributed to RIS-based communication systems, delivering robust performance across diverse environmental conditions and enabling low-complexity inference. Various deep learning approaches have been proposed to improve cascaded channel estimation for massive MIMO systems, significantly improving overall system performance [DL5, DL1, DL2, DL3, DL4, DL6]. However, these techniques face scalability challenges and are typically limited to massive MIMO systems, where only tens of antennas and BS elements are considered. Extending the training setup to RIS-enabled XL-MIMO systems, involving hundreds of RIS elements and thousands of antennas as proposed in these models, would not be a viable solution due to various reasons. Firstly, these techniques use tens of thousands of samples to train their models. Processing and storing such large datasets in RIS-assisted XL-MIMO-based systems can lead to inefficiencies, introduce substantial computational and memory demands, and increase both training time and resource consumption [DL1]. Furthermore, training these models on large datasets requires substantial energy resources, making them not only energy inefficient but also raising environmental concerns due to increasing energy consumption. Additionally, the need for highly specialized high-performance hardware to handle the training of such expensive systems adds considerable costs, making it financially challenging to scale these solutions. Finally, with edge computing gaining attention in recent years as a key enabler for distributed low-latency applications, the deployment of these complex models in edge environments becomes a significant obstacle [FL]. The limited computational resources and energy constraints at the edge create a major barrier, hindering the practical implementation of these techniques in future communication networks.
@@
Addressing the aforementioned challenges requires a compact and efficient representation of the cascaded channel. We, therefore, exploit the spatial correlation present in the cascaded channel, as discussed in [sp1, sp2, sp3], to enable computationally efficient model training suitable for edge devices in resource-limited settings. To take advantage of this property, we introduce a patch-based training framework that leverages the inherent spatial correlations in the cascaded channel. In the proposed framework, small, localized patches are extracted from each high-dimensional data point to effectively capture salient features necessary for accurate channel estimation. This patch-wise processing significantly reduces input dimensionality and overall system complexity, thereby enabling scalable and computationally efficient training, particularly suitable for XL-MIMO systems deployed at the edge. Furthermore, employing efficient deep learning techniques, such as encoder-decoder architecture, enables the transformation of representative patches into compact latent representations, thereby facilitating more effective learning of spatial representation [UNET1]. These latent features effectively preserve and highlight spatial dependencies between RIS elements and BS antennas, facilitating the learning of inherent correlations while substantially reducing training complexity. Furthermore, leveraging the encoder-decoder architecture also helps break the correlation of noise in the data, effectively isolating meaningful patterns from unwanted distortions [AE1]. For example, in [GM2], the model demonstrated strong potential for channel estimation in MIMO systems.
@@
Therefore, inspired by the encoder–decoder architecture, our proposed channel estimation framework incorporates multi-scale feature extraction and aggregation mechanisms, allowing for more effective representation of salient structures within noisy inputs, thereby significantly improving estimation performance. This approach captures spatial features efficiently and enhances estimation accuracy, particularly in resource-constrained environments such as edge devices. To the best of our knowledge, this is the first work to propose a deep learning-based solution for channel estimation in RIS-assisted XL-MIMO systems. Moreover, the channel estimation is formulated as a denoising problem aimed at recovering clean channel information from noisy observations. The proposed method significantly improves the accuracy of the estimation and reduces computational complexity, regardless of the increasing number of antennas and RIS elements in XL-MIMO systems. Extensive simulations under diverse environmental conditions validate the scalability and effectiveness of the proposed framework, demonstrating substantial gains in cascaded channel estimation and system efficiency for RIS-enabled XL-MIMO deployments.
@@ I-A Literature Review
Efficient cascaded channel estimation has attracted significant attention in recent years due to the promising benefits of RIS. Existing research efforts can be broadly categorized into two main approaches: traditional model-driven methods and data-driven techniques. Traditional model-based approaches for cascaded channel estimation remain widely used due to their ability to operate without training data, allowing direct deployment without prior model training. In [OP1], the authors tackled the cascaded channel estimation problem by exploiting the sparsity and structural correlations in mmWave environments. Their primary optimizations focused on adjusting the phase shifts of RIS elements and minimizing the pilot overhead required for successive channel updates. Moreover, [BLMMSE] proposes a binary reflection-based channel estimation method, activating one element at a time while the others remain off, with channel estimates obtained via Least Squares (LS).
@@
Furthermore, in [OP2], the authors formulated the cascaded channel estimation task using an alternating optimization algorithm combined with a covariance method, which they solved through a successive convex approximation algorithm. Their findings demonstrated enhanced channel estimation accuracy while reducing pilot overhead. The authors in [OP3] proposed a channel estimation protocol that leveraged compressive sensing techniques to reduce training overhead and estimate the cascaded channel as a sparse recovery problem. They introduced a two-step joint channel estimation framework to further improve estimation performance by exploiting the block-sparsity structure shared among multiple users. Despite the adequate performance of these approaches  [OP1, OP2, OP3], they rely on numerous iterative steps to achieve optimal results, which limits their efficiency in real-time, low-latency applications when applied to XL-MIMO systems involving hundreds of RIS elements and thousands of antennas. As a result, model-based approaches become impractical for such large-scale deployments due to their excessive complexity and resource demands.
@@
The pioneering work that introduced deep learning for cascaded channel estimation was presented in [DL5]. The proposed framework consisted of two convolutional neural networks (CNNs), each comprising nine layers, designed to estimate both the direct channel between the BS and the user and the cascaded channel between the BS and the user through an RIS. The simulation results validated the effectiveness of the proposed model, demonstrating improvements over traditional methods such as LS and multilayer perceptrons. Moreover, in [DL1], the authors proposed a deep residual network (CDRN) method using a CNN-based architecture to effectively remove residual noise from noisy pilot signals, ultimately enhancing estimation accuracy. Additionally, [DL2] introduced a bidirectional long short-term memory (Bi-LSTM) approach for channel estimation. Bi-LSTM processed past and future information simultaneously through two LSTM layers, one forward and one backward, thereby enhancing feature extraction and achieving adequate accuracy in channel estimation.
@@
Similarly, the work in [DL3] interprets the received pilot signal as a low-resolution image and formulates channel estimation as a super-resolution reconstruction problem. This approach employed a CNN to extract detailed features, followed by a denoising CNN to eliminate additive noise. Another study in [DL6] proposed a model-driven deep unfolding approach for cascaded channel estimation. The proposed scheme leverages the rank deficiency of the cascaded channel to achieve accurate estimation while significantly reducing communication overhead. In [DL4], a self-supervised learning framework was developed for channel estimation, designed to operate under both supervised and unsupervised conditions. Specifically, in the absence of ground-truth data, the method synthesized input samples by introducing controlled noise into the received signal, using the resulting noisy observations as the corresponding labels. Although this approach yielded notable results, it faced limitations in real-world scenarios where estimates of accurate noise variance are often unavailable. Although the techniques mentioned above [DL5, DL1, DL2, DL3, DL4, DL6] demonstrate satisfactory performance, they still encounter challenges related to scalability and estimation accuracy and rely on unrealistic dataset assumptions that may not adequately reflect real-world scenarios.
@@
A significant limitation of data-driven techniques, as mentioned earlier, is that most proposed schemes are trained and evaluated on small-scale systems, typically involving only tens of RIS and antenna elements. Although these schemes perform effectively in moderate-scale scenarios, their scalability becomes a critical bottleneck when applied to large-scale systems comprising hundreds of RIS elements and thousands of antennas, rendering them impractical for RIS-enabled XL-MIMO deployments [RIS]. Training such large systems requires complex, high-end computational infrastructure to meet the associated processing demands, posing a major challenge to the feasibility of these approaches. Moreover, several studies [RIS, OP1, OP3] have highlighted the presence of spatial correlation in the cascaded channel. In particular, some works have conducted detailed analyses of spatially correlated RIS and antenna elements, providing valuable insights into the impact of correlation on system performance [sp1, sp2, sp3]. Motivated by these observations, this paper leverages the spatial correlation inherent in the cascaded channel to reduce model complexity, thereby enabling efficient deployment of XL-MIMO systems on resource-constrained devices. To ensure effective training, a multi-scale feature extraction and aggregation architecture is adopted.